# A Simple Unified Uncertainty-Guided Framework for Offline-to-Online Reinforcement Learning

### v2

@@ 5.2 Comparisons on D4RL Benchmarks
Table 1: Comparison of the averaged D4RL score on MuJoCo tasks with TD3+BC as the offline RL backbone method. r = random, m = medium, m-r = medium-replay. We report the mean and standard deviation over 5 seeds.
TD3+BC offline-ft online-ft BR O3F APL PEX SUNG
@@
Total 389.6 454.9 532.9 590.9 637.7 647.1 521.2 705.7
Results on MuJoCo. Results for MuJoCo domains with TD3+BC and CQL as backbone offline RL method are shown in Table 1 and 2, respectively. We find that additional online interactions can bring further performance improvement in most settings. However, two exceptions are O3F and PEX, when combined with CQL. As pointed in Mark et al. (2022), O3F is not compatible with value regularization based offline RL methods. This is also observed by Nakamoto et al. (2023) and we provide detailed explanations in Appendix B. In contrast, PEX trains a new policy from scratch, and cannot fully benefit from reusing good initialization from pretraining, thereby leading to limited performance.
Results on AntMaze. We also provide results for the more challenging AntMaze domains with SPOT and CQL as backbone offline RL method. We defer the results and detailed analyses to Appendix D.1. We observe that SUNG outperforms all the baselines in most settings.
5.3 Ablation Studies
In this subsection, we perform an ablation study over the components in SUNG in Fig. 2. Refer to Appendix D.4 for implementation details and extra ablation studies.
Table 2: Comparison of the averaged D4RL score on MuJoCo tasks with CQL as the offline RL backbone method. r = random, m = medium, m-r = medium-replay. We report the mean and standard deviation over 5 seeds. Note that O3F is omitted due to its divergence in this setting.
CQL offline-ft online-ft BR APL PEX SUNG
@@ 5.3 Ablation Studies
Total 466.9 501.1 556.4 558.0 617.8 371.5 683.4
Ablation on Optimistic Exploration. First, we evaluate (a) SUNG without optimistic exploration strategy, and find that the removal brings significant negative effects when combined with TD3+BC, but no remarkable effects when combined with CQL. The reason for this discrepancy is that CQL is built on top of SAC, which can gradually recover exploration power through max-entropy RL framework during online finetuning. Then, we ablate (b) uncertainty and (c) Q value in the optimistic exploration strategy to derive two variant strategies - one greedy for Q value and another greedy for uncertainty. As expected, both ablations degrade the performance to varying degrees, underscoring the significance of both Q value and uncertainty for exploration. Notably, the removal of uncertainty in SUNG when combined with CQL causes significant performance degradation, which can be attributed to similar reason detailed in Appendix B.
Ablation on Adaptive Exploitation. We evaluate (d) SUNG without the adaptive exploitation method, and observe that performance for finetuning TD3+BC and CQL significantly deteriorates on most tasks. This emphasizes the necessity of addressing the state-action distribution shift issue.
Figure 2: Performance difference of an ablation study of SUNG combined with TD3+BC and CQL, compared with the full algorithm. Opt. = Optimistic, Unc. = Uncertainty, Adp. = Adaptive. hc = HalfCheetah, hop = Hopper, w = Walker, r = random, m = medium, mr = medium-replay.
Figure 3: Comparing performance with different order of ranking criteria and finalist action set size $k$ used in optimistic exploration.
Figure 3: Comparing performance with different order of ranking criteria and finalist action set size $k$ used in optimistic exploration.
Figure 4: Comparing performance with varying percent $p$ of identified OOD samples from the mini-batch used in adaptive exploitation.
Ablation on Uncertainty Quantification. Finally, we (e) replace the VAE-based density estimator with the standard deviation of double Q values for uncertainty quantification Ciosek et al. (2019). As expected, the results show that such replacement leads to performance deterioration on most tasks, since it is not sufficient to provide reliable uncertainty quantification. While utilizing the ensemble technique can eliminate this issue Lee et al. (2021); Chen et al. (2017), it can significantly increase computational costs. We leave the exploration of alternative methods as future work.
5.4 Hyper-parameter Analysis
Analysis of Optimistic Exploration. First, we investigate the order of the ranking criteria and the finalist action set size $k$ in optimistic exploration. The aggregated results averaged over MuJoCo domains are shown in Fig. 4. As expected in Appendix B, SUNG with TD3+BC can accommodate preference for both high uncertainty and high value, whereas SUNG with CQL can only accommodate preference for high uncertainty. Hence, we choose the order with Q first for TD3+BC, and uncertainty first for CQL. Besides, we remark that SUNG consistently outperforms the best baseline with any choice of $k$, which further verifies the superiority of SUNG.
Analysis of Adaptive Exploitation. Moreover, we explore the choice of percent $p$ of identified OOD samples from the mini-batch in adaptive exploitation. The aggregated results averaged over MuJoCo domains are shown in Fig. 4. Predictably, setting $p$ too conservatively or too aggressively may adversely degrade the performance. We find that $p=5$ performs the best when combined with either TD3+BC or CQL, achieving the desired trade-off between performance and stability.
6 Conclusion, Limitations, and Broader Impacts
This paper studies how to efficiently finetune pretrained offline RL agents via online interactions. We present SUNG, a simple unified uncertainty-guided framework, to tackle both challenges of constrained exploratory behavior and state-action distribution shift. SUNG leverages the quantified uncertainty to guide optimistic exploration and adaptive exploitation. Empirical results across different backbone offline RL methods, environments and datasets verify the superiority of SUNG.
In terms of limitations, the proposed offline-to-online framework cannot be directly applied to recent advances in offline RL, such as reward-conditioned RL based methods Chen et al. (2021); Emmons et al. (2022) and on-policy methods Zhuang et al. (2023). However, we argue that the insights of SUNG are generally applicable in principle. We leave this as future work. Regarding societal impacts, we do not anticipate any negative consequences from using our framework in practice. We provide more detailed discussions in Appendix E.
Appendix
Appendix A Implementation Details of SUNG
In this section, we provide two implementation cases of SUNG: one bridges TD3+BC Fujimoto and Gu [2021] and TD3 Fujimoto et al. [2018], and another bridges CQL Kumar et al. [2020] and SAC Haarnoja et al. [2018]. Note that existing offline RL methods are typically built on top of online off-policy RL methods. Thus, we utilize the corresponding online RL objectives for online finetuning to keep consistent.
A.1 SUNG (TD3+BC): Bridging TD3+BC and TD3
First, we show how SUNG (TD3+BC) bridges TD3+BC and TD3 to finetune the deterministic policy $\pi_{\phi}(s)$ and two Q value functions $Q_{\theta_{1}}(s,a),Q_{\theta_{2}}(s,a)$.
In terms of exploration, the behavior policy of SUNG (TD3+BC) is compatible with preference for either high Q value or high uncertainty, as demonstrated in the main body. Here, we take preference for high Q value as example. As TD3 utilizes the deterministic policy, we first generate a candidate action set $\mathcal{C}=\{a_{i}=\pi_{\phi}(s)+\epsilon_{i}\}_{i=1}^{N}$, where $\epsilon_{i}\sim\mathcal{N}(0,\delta)$ is the Gaussian noise. Then, candidate actions with top-$k$ Q values are selected as finalist action set $\mathcal{C}_{f}$. Finally, we sample the action for interaction from a categorical distribution $\mathrm{P}(a_{i}),a_{i}\in\mathcal{C}_{f}$ in Eq. (8) according to the uncertainty. In practice, we set $N=100$ and $k=10$ to interact with the environment by selecting actions that possess high value and relatively high uncertainty.
In terms of exploitation, SUNG (TD3+BC) follows TD3+BC and TD3 to take the standard TD learning for policy evaluation, and adopts the proposed adaptive exploitation method for policy improvement as
$\mathcal{L}_{\pi}^{\mathrm{adp,TD3+BC}}(\phi)=\mathbb{E}_{s\sim\mathcal{B}_{\mathrm{OORB}}}\left[-Q_{\theta}\left(s,\pi_{\phi}\left(s\right)\right)\right]+\lambda\mathbb{E}_{(s,a)\sim\mathcal{B}_{\mathrm{OORB}}}\left[\mathcal{I}\left(s,\pi_{\phi}\left(s\right)\right)\left(\pi_{\phi}\left(s\right)-a\right)^{2}\right].$ (11)
Integrating the details above, we give a detailed algorithm description of SUNG (TD3+BC) in Algorithm 2.
Algorithm 2 SUNG (TD3+BC): Bridging TD3+BC and TD3
1:  Input: Dataset $\mathcal{D}=\left\{\left(s,a,r,s^{\prime}\right)\right\}$
2:  // Offline Pretraining
3:  Initialize Q network $Q_{\theta}$, policy network $\pi_{\phi}$, VAE network with $p_{\psi}$ and $q_{\varphi}$
4:  for $t=1$ to $T_{1}$ do
5:     Sample minibatch of transitions $\left(s,a\right)\sim\mathcal{D}$
6:     Update $\theta$ and $\phi$ minimizing $\mathcal{L}_{Q}(\theta)$ in Eq. (1) and $\mathcal{L}^{\mathrm{TD3+BC}}_{\pi}(\phi)$ in Eq. (3)
7:     Update $\psi,\varphi$ minimizing $\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)$ in Eq. (6)
8:  end for
9:  // Online Finetuning
10:  Initialize OORB with $\mathcal{B}_{\mathrm{OORB}}\leftarrow\{\mathcal{B}=\emptyset,\mathcal{D}\}$
11:  for $t=1$ to $T_{2}$ do
12:     // Optimistic Exploration
13:     Generate candidate action set $\mathcal{C}=\{a_{i}\}_{i=1}^{N}$, where $a_{i}=\pi_{\phi}(s_{t})+\epsilon_{i}$, $\epsilon_{i}\sim\mathcal{N}(0,\delta)$
14:     Select actions with top-$k$ Q-values as finalist action set $\mathcal{C}_{f}$
15:     Sample the action $a_{t}\sim\mathrm{P}(a_{i}),a_{i}\in\mathcal{C}_{f}$ in Eq. (8) according to the uncertainty
16:     Execute action $a_{t}$, collect $s_{t+1}$ and ${r_{t}}$ from the environment
17:     Store transitions $(s_{t},a_{t},r_{t},s_{t+1})$ into $\mathcal{B}$ and $\mathcal{D}$
18:     // Adaptive Exploitation
19:     Sample minibatch of transitions $\left(s,a,r,s^{\prime}\right)\sim\mathcal{B}_{\mathrm{OORB}}$
20:     Construct the OOD identifier $\mathcal{I}$ in Eq. (10)
21:     Update $\theta$ minimizing $\mathcal{L}_{Q}(\theta)$ in Eq. (1)
22:     Update $\phi$ minimizing $\mathcal{L}_{\pi}^{\mathrm{adp,TD3+BC}}(\phi)$ in Eq. (11)
23:     Update $\psi,\varphi$ minimizing $\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)$ in Eq. (6)
24:  end for
A.2 SUNG (CQL): Bridging CQL and SAC
Then, we show how SUNG (CQL) bridges CQL and SAC to finetune the stochastic policy $\pi_{\phi}(\cdot|s)$ and two Q value functions $Q_{\theta_{1}}(s,a),Q_{\theta_{2}}(s,a)$.
In terms of exploration, the behavior policy of SUNG (CQL) is only compatible with preference for high uncertainty, as demonstrated in the main body. As SAC utilizes the stochastic policy, we first generate a candidate action set $\mathcal{C}=\{a_{i}\sim\pi_{\phi}(\cdot|s)\}_{i=1}^{N}$. Then, candidate actions with top-$k$ uncertainty are selected as finalist action set $\mathcal{C}_{f}$. Finally, we sample the action for interaction from a categorical distribution $\mathrm{P}(a_{i}),a_{i}\in\mathcal{C}_{f}$ similar to Eq. (8) according to the Q value. In practice, we set $N=100$ and $k=20$. Note that we slightly increase the value of $k$ to avoid the selected actions with too high uncertainty for sample efficiency. As such, we can interact with the environment by selecting actions that possess relatively high uncertainty and relatively high Q value.
In terms of exploitation, SUNG (CQL) follows CQL and SAC to perform policy improvement as
$\mathcal{L}_{\pi}^{\mathrm{SAC}}(\phi)=\mathbb{E}_{s\sim\mathcal{B}_{\mathrm{OORB}},a\sim\pi_{\phi}(\cdot|s)}\left[Q_{\theta}(s,a)-\log\pi_{\phi}(a|s)\right],$ (12)
and adopts the proposed adaptive exploitation method for policy evaluation as
$\mathcal{L}_{Q}^{\mathrm{adp,CQL}}(\theta)$ $=\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{B}_{\mathrm{OORB}}}\left[\left(Q_{\theta}(s,a)-r-\gamma Q_{\bar{\theta}}\left(s^{\prime},\pi_{\phi}\left(s^{\prime}\right)\right)\right)^{2}\right]$ (13)
$+\lambda\mathbb{E}_{s\sim\mathcal{B}_{\mathrm{OORB}}}\left[\mathcal{I}\left(s,\pi_{\phi}\left(s\right)\right)\left(\mathbb{E}_{a\sim\pi_{\phi}}\left[Q_{\theta}\left(s,a\right)\right]-\mathbb{E}_{a\sim\mathcal{B}_{\mathrm{OORB}}}\left[Q_{\theta}\left(s,a\right)\right]\right)\right],$
where $\pi_{\phi}(s)$ denotes the mean of the Gaussian stochastic policy. Integrating the details above, we give a detailed algorithm description of SUNG (CQL) in Algorithm 3.
Algorithm 3 SUNG (CQL): Bridging CQL and SAC
1:  Input: Dataset $\mathcal{D}=\left\{\left(s,a,r,s^{\prime}\right)\right\}$
2:  // Offline Pretraining
3:  Initialize Q network $Q_{\theta}$, policy network $\pi_{\phi}$, VAE network with $p_{\psi}$ and $q_{\varphi}$
4:  for $t=1$ to $T_{1}$ do
5:     Sample minibatch of transitions $\left(s,a\right)\sim\mathcal{D}$
6:     Update $\theta$ and $\phi$ minimizing $\mathcal{L}_{Q}^{\mathrm{CQL}}(\theta)$ in Eq. (4) and $\mathcal{L}_{\pi}^{\mathrm{SAC}}(\phi)$ in Eq. (12)
7:     Update $\psi,\varphi$ minimizing $\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)$ in Eq. (6)
8:  end for
9:  // Online Finetuning
10:  Initialize OORB with $\mathcal{B}_{\mathrm{OORB}}\leftarrow\{\mathcal{B}=\emptyset,\mathcal{D}\}$
11:  for $t=1$ to $T_{2}$ do
12:     // Optimistic Exploration
13:     Generate candidate action set $\mathcal{C}=\{a_{i}\sim\pi_{\phi}(\cdot|s)\}_{i=1}^{N}$
14:     Select actions with top-$k$ uncertainty as finalist action set $\mathcal{C}_{f}$
15:     Sample the action $a_{t}\sim\mathrm{P}(a_{i}),a_{i}\in\mathcal{C}_{f}$ similar to Eq. (8) according to the Q value
16:     Execute action $a_{t}$, collect $s_{t+1}$ and ${r_{t}}$ from the environment
17:     Store transitions $(s_{t},a_{t},r_{t},s_{t+1})$ into $\mathcal{B}$ and $\mathcal{D}$
18:     // Adaptive Exploitation
19:     Sample minibatch of transitions $\left(s,a,r,s^{\prime}\right)\sim\mathcal{B}_{\mathrm{OORB}}$
20:     Construct the OOD identifier $\mathcal{I}$ in Eq. (10)
21:     Update $\theta$ minimizing $\mathcal{L}_{Q}^{\mathrm{adp,CQL}}(\theta)$ in Eq. (13)
22:     Update $\phi$ minimizing $\mathcal{L}_{\pi}^{\mathrm{SAC}}(\phi)$ in Eq. (12)
23:     Update $\psi,\varphi$ minimizing $\mathcal{L}_{\mathrm{ELBO}}(s,a;\psi,\varphi)$ in Eq. (6)
24:  end for
Appendix B Why Exploration with Preference for High Value Fails When Finetuning CQL?
In empirical results, we find that O3F Mark et al. [2022], which selects actions with preference for high Q value during exploration, fails when combined with CQL Kumar et al. [2020]. Besides, SUNG with preference of high Q value for exploration also fails when combined with CQL. Now, we give a qualitative analysis of this reason. To begin with, we recall the practical implementation of the policy evaluation in CQL:
$\arg\min_{Q}\mathbb{E}_{s\in\mathcal{D}}\left[\log\sum_{a}\exp\left(Q(s,a)\right)-\mathbb{E}_{a\sim\pi_{\beta}(a|s)}\left[Q(s,a)\right]\right]+\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[(Q(s,a)-y)^{2}\right],$ (14)
where $\pi_{\beta}$ denotes the behavior policy of the offline dataset and $y$ denotes the standard TD target for policy evaluation.
There are three terms in the above equation: given a state, the first term minimizes Q values of all the actions; the second term maximizes Q values of the action collected in the dataset; the third term is the standard TD learning. As such, given state $s$, when we choose high-value action $a_{h}$ for exploration, the transition $(s,a_{h},s^{\prime},r)$ will be collected into the replay buffer. Then, during exploitation, the combination of the first and second term in the equation leads to a situation where most unseen actions have low value and will have lower value due to the first term, while few seen actions have high value and will have higher value due to the second term. As the third term involves bootstrapping, the high-value action will be updated to have a lower value, leading to a consistent decrease in the values of all actions with each gradient step. This can ultimately cause the collapse of the Q value functions, further resulting in poor policy through propagation in the policy improvement.
We remark that SUNG can solve this issue by altering the ranking criteria for the finalist action set and $k$ to show preference for high uncertainty while relatively high Q value during exploration.
Appendix C More Experimental Details
C.1 Experimental Setting Details
For MuJoCo tasks, we follow prior work Zhang et al. [2023] to omit the evaluation on medium-expert and expert datasets, where offline pre-trained agents can already achieve expert-level performance and do not need further finetuning. All evaluations use datasets of the latest released "-v2" version. All the experiments are repeated with 5 different random seeds. For each evaluation, we run 10 episodes for MuJoCo tasks and 100 episodes for AntMaze tasks. We report the average D4RL score over last 3 evaluations for MuJoCo tasks, and last 10 evaluations for AntMaze tasks.
For offline pretraining, we first train the agents with 1M gradient steps by taking the suggested standard pipeline of TD3+BC and CQL:
Then, the last checkpoint of the pre-trained agents are used to warm start the online finetuning. Since the authors of SPOT Wu et al. [2022] have provided the checkpoints of the pre-trained agents in https://github.com/thuml/SPOT, we directly adopt these for online finetuning.
We conduct all the experiments on 2 NIVIDA 3090 GPUs and Intel Xeon Gold 6230 CPU at 2.10GHz.
C.2 Implementation Details and Hyper-parameters
For VAE training, we follow the official implementation of BCQ Fujimoto et al. [2019] and SPOT Wu et al. [2022]. Following SPOT, we normalize the states in the dataset for MuJoCo domains but do not normalize the states for AntMaze domains. Different from SPOT that utilizes VAE to estimate the density of the behavior policy, SUNG utilizes VAE to estimate the state-action visitation density. Thus, we set the latent dimension as $2\times(\mathrm{state~{}dim}+\mathrm{action~{}dim})$ instead of $2\times\mathrm{action~{}dim}$. Detailed hyper-parameters for VAE training are listed in Table 3. Note that we simply take the hyper-parameters reported in SPOT Wu et al. [2022] and omit hyper-parameter tuning for VAE training. Accordingly, further performance improvement is expected from careful tuning.
For online finetuning, we follow the same standard architecture and the same hyper-parameter of TD3+BC, CQL and SPOT for all the baselines. Note that we do not tune any hyper-parameter of backbone RL and architecture in SUNG for fair comparison. In terms of baseline reproduction, we strictly follow their official implementation and hyper-parameters reported in their original paper to finetune agents pre-trained with TD3+BC, CQL and SPOT for MuJoCo and AntMaze domains. One exception is that the heavy ensemble technique is removed in BR for fair comparison. Note that we also equip the baseline with OORB for fair comparison. All the detailed hyper-parameters for online finetuning are presented in Table 4.
Table 3: Hyper-parameters of VAE for uncertainty quantification in SUNG.
Hyper-parameter Value
VAE training Optimizer Adam
Learning rate 1e-3
Batch size 256
Gradient steps 1e5
KL term weight 0.5
State Normalization True for MuJoCo
False for AntMaze
VAE architecture Encoder hidden num 750
Encoder layers 3
Latent dim 2 $\times$ (state dim + action dim)
Decoder hidden dim 750
Decoder layers 3
Table 4: Hyper-parameters of online finetuning in SUNG. Note that we do not tune any hyper-parameter in backbone RL and architecture.
Hyper-parameter MuJoCo AntMaze
TD3+BC CQL SPOT CQL
Backbone RL Optimizer Adam Adam Adam Adam
Actor learning rate 3e-4 1e-4 1e-4 1e-4
Critic learning rate 3e-4 3e-4 3e-4 3e-4
Batch size 256 256 256 256
Discount factor 0.99 0.99 0.99 0.99
Target update rate 5e-3 5e-3 5e-3 5e-3
Architecture Actor hidden num 256 256 256 256
Actor layers 3 2 3 2
Critic hidden num 256 256 256 256
Actor layers 3 2 3 3
SUNG Candidate action num $N$ 100 100 100 100
Finalist action num $k$ 10 20 10 20
Order of ranking criteria qu uq qu uq
Softmax temperature $\alpha$ 1.0 1.0 1.0 1.0
OOD sample percentage $p$ 5 10 99 90
OORB Online buffer size 2e4 2e4 2e4 2e4
Offline buffer size 2e6 2e6 2e6 2e6
OORB probability $p_{\mathrm{OORB}}$ 0.1 0.1 0.2 0.7
Appendix D More Experimental Results
D.1 Experiments on AntMaze
Table 5: Comparison of the averaged D4RL score on AntMaze tasks with SPOT as the offline RL backbone method. We report the mean and standard deviation over 5 seeds.
SPOT offline-ft online-ft BR O3F APL PEX SUNG
@@ D.1 Experiments on AntMaze
3.0
$\pm$0.9
3.8
$\pm$0.9
0.0
$\pm$0.0
19.8 $\pm$1.6
Total 367.8 344.6 88.8 99.3 369.7 362.8 12.3 414.5

### v3

@@ V-A4 Implementation details and hyper-parameters
TABLE II: Comparison of the averaged D4RL score on MuJoCo tasks with TD3+BC as the offline RL backbone method. We report the mean and standard deviation over 5 seeds.
TD3+BC offline-ft online-ft BR O3F APL PEX PROTO SUNG
@@
Total 389.6 454.9 532.9 590.9 637.7 647.1 521.2 530.2 705.7
Avg Improvement / 16.76% 36.78% 51.66% 63.68% 66.09% 33.77% 36.08% 81.13%
TABLE III: Comparison of the averaged D4RL score on MuJoCo tasks with CQL as the offline RL backbone method. We report the mean and standard deviation over 5 seeds. Note that O3F is omitted due to its divergence in this setting.
CQL offline-ft online-ft BR APL PEX PROTO SUNG
@@
Total 466.9 501.1 556.4 558.0 617.8 371.5 536.1 683.4
Avg Improvement / 7.32% 19.16% 19.51% 32.31% -20.43% 14.82% 46.36%
V-A5 Baselines
We compare SUNG with the following baselines for online finetuning:
  • offline-ft leverages online interactions by performing offline RL objectives.
  • online-ft leverages online interactions by performing online RL objectives that correspond to the ones used for pretraining.
  • BR [6] prioritizes near-on-policy transitions from the replay buffer.
  • O3F [11] utilizes knowledge contained in value functions to guide exploration.
  • APL [12] performs adaptive policy learning by incorporating different advantages of offline and online data.
  • PEX [7] trains a new policy from scratch by utilizing the pre-trained policy for exploration.
  • PROTO [39] introduces an iterative policy regularization scheme to achieve stable finetuning performance.
Note that we also include IQL [32], ODT [8], and ACA [33] for comparison with a separate subsection in Section V-C.
V-B Comparisons on D4RL Benchmarks (RQ1)
In this subsection, we perform the overall performance comparison with different offline-to-online RL methods on both MuJoCo and AntMaze domains in D4RL benchmark. The full learning curves for all the settings are provided in the supplementary material.
V-B1 Results on MuJoCo
Results for MuJoCo domains with TD3+BC and CQL as backbone offline RL method are shown in Table II and III, respectively. We find that additional online interactions can bring further performance improvement in most settings. However, two exceptions are O3F and PEX, when combined with CQL. As pointed in [11], O3F is not compatible with value regularization based offline RL methods. In contrast, PEX trains a new policy from scratch, and cannot fully benefit from reusing good initialization from pretraining, thereby leading to limited performance.
TABLE IV: Comparison of the averaged D4RL score on AntMaze tasks with SPOT and CQL as the offline RL backbone method. We report the mean and standard deviation over 5 seeds.
SPOT offline-ft O3F APL SUNG CQL offline-ft O3F APL SUNG
@@ V-B1 Results on MuJoCo
44.1 $\pm$12.3
Total 367.8 344.6 369.7 362.8 414.5 319.6 360.4 246.7 155.6 416.1
Avg Improvement / -6.30% 0.51% -1.35% 12.69% / 12.76% -22.80% -51.31% 30.19%
V-B2 Results on AntMaze
We conduct experiments on more challenging AntMaze domains, which require the agents to deal with the sparse rewards and stitch fragments of sub-optimal trajectories. Results for AntMaze domains with SPOT and CQL as backbone offline RL method are shown in Table IV. We omit the results of online-ft, BR, PEX, and PROTO due to their poor performance. This stems from the need for explicit or implicit behavior cloning that is offered by offline RL objectives, which is required to handle the sparse rewards and stitch sub-optimal sub-trajectories. However, online-ft, BR, PEX, and PROTO are developed on top of purely online RL methods, thus failing in most of the tasks.
Furthermore, we highlight that SUNG exhibits an additional 12.43% and 17.43% offline-to-online improvement over the best-performing baseline when combined with SPOT and CQL, respectively. This demonstrates the superiority of SUNG for discovering informative state-action pairs that can lead to high outcomes, thereby providing valuable behavior patterns for agents to learn from. Moreover, SUNG achieves flexible trade-off between offline and online RL objectives by tuning $p$, which enables stitching sub-optimal sub-trajectories and overcoming the sparse reward issue.
In the following subsections, we mainly focus on analyses in MuJoCo domains, given their wide adoption in previous works [36, 6, 12].
V-C Comparison with Other Offline-to-Online RL Methods (RQ1)
In this subsection, we compare SUNG with other offline-to-online RL methods that are designed for one specific offline RL method. Concretely, we select the following baselines: IQL [32], ODT [8] and ACA [33]. We follow the original experimental setting in ODT to take 200K environment steps and only consider tasks with medium and medium-replay offline datasets for online finetuning to achieve fair comparison. The results are directly copied from [33].
We present experimental results of the final performance of different offline-to-online RL methods with 200K environment steps. As shown in Fig. 3, for almost all the settings, SUNG when combined with either TD3+BC or CQL outperforms other methods by a large margin in terms of final performance. IQL presents conservative online finetuning, which can stably improve the performance but derive limited performance under a limited interaction setting, i.e., 200K environment steps. ODT learns a reward-conditioned policy in a supervised learning paradigm, thereby struggling to fully benefit from the trail-and-error paradigm used in traditional online RL. However, ODT enables users to customize the behavior of the agents by specifying different conditions, which is not featured by traditional RL.
Figure 3: Final performance of different offline-to-online RL methods with 200K environment steps. We report the mean D4RL score across 5 random seeds. hc = halfcheetah, hop = hopper, w = walker2d, m = medium, mr = medium-replay.
V-D Ablation Studies (RQ2)
In this subsection, we showcase experimental results in Fig. 4 to validate the effectiveness of each component of SUNG.
Figure 4: Performance difference of an ablation study of SUNG combined with TD3+BC and CQL, compared with the full algorithm. We report the mean performance difference across 5 different random seeds. Opt. = Optimistic, Unc. = Uncertainty, Adp. = Adaptive. hc = halfcheetah, hop = hopper, w = walker2d, r = random, m = medium, mr = medium-replay.
V-D1 Ablation on Optimistic Exploration
We evaluate (a) SUNG w/o Optimistic Exploration, i.e., we replace the proposed optimistic exploration strategy by the default exploration provided by TD3 and SAC. For TD3, we select action with Gaussian noise for exploration, i.e., $\pi_{E}(s)=\pi_{\phi}(s)+\epsilon,\epsilon\sim\mathcal{N}(0,\delta)$. For SAC, we sample action from the policy for exploration, i.e., $\pi_{E}(s)=a,a\sim\pi_{\phi}(\cdot|s)$. We find that the removal brings significant negative effects when combined with TD3+BC, but no remarkable effects when combined with CQL. The reason for this discrepancy is that CQL is built on top of SAC, which can gradually recover exploration power through max-entropy RL framework during online finetuning.
Moreover, we ablate two key components in the optimistic exploration strategy, uncertainty and Q value, to derive two variant strategies: (b) Optimistic Exploration w/o Uncertainty and (c) Optimistic Exploration w/o Q. For the former, we remove the uncertainty from the proposed exploration strategy to greedily select action that maximizes Q value, i.e., $\pi_{E}(s)=\arg\max_{a\in\mathcal{C}}Q(s,a)$. For the latter, we remove the Q value from the proposed exploration strategy to greedily select action that maximizes the uncertainty, i.e., $\pi_{E}(s)=\arg\max_{a\in\mathcal{C}}\mathcal{U}(s,a)$. As expected, both ablations degrade the performance to varying degrees, underscoring the significance of both Q value and uncertainty for exploration. Notably, the removal of uncertainty in SUNG when combined with CQL causes significant performance degradation, which can be attributed to similar reason detailed in [11].
V-D2 Ablation on Adaptive Exploitation
We evaluate (d) SUNG w/o Adaptive Exploitation, which replaces the proposed adaptive exploitation method by the standard online RL objective of TD3 and SAC, respectively. From the corresponding experimental results, we note a significant degradation in performance for finetuning TD3+BC and CQL across most tasks. This emphasizes the necessity of addressing the state-action distribution shift issue.
V-D3 Ablation on Uncertainty Quantification
We introduce the variant: (e) Uncertainty Estimation w/ Q std, which replaces the VAE-based density estimator with the standard deviation of double Q values for uncertainty quantification, i.e., $\mathcal{U}(s,a)=\sigma(Q_{\theta_{1}}(s,a),Q_{\theta_{2}}(s,a))=|Q_{\theta_{1}}(s,a)-Q_{\theta_{2}}(s,a)|$. As expected, the results show that such replacement leads to performance deterioration on most tasks, since it is not sufficient to provide reliable uncertainty quantification. While utilizing the ensemble technique can eliminate this issue [13, 54], it can significantly increase computational costs. We leave the exploration of alternative methods as future work.
Figure 5: Comparing performance on MuJoCo domains with different finalist action set size $k$ used in the optimistic exploration strategy.
Figure 6: Comparing performance on MuJoCo domains with varying percent $p$ of identified OOD samples from the mini-batch used in the adaptive exploitation method.
V-E Hyper-parameter Analysis (RQ3)
In this subsection, we give a hyper-parameter analysis of SUNG, focusing on two important hyperparameters: final action set size $k$ and identified OOD sample percentage $p$.
V-E1 Analysis of Finalist Action Set Size $k$
First, we investigate the finalist action set size $k$ in optimistic exploration. We tune it in the range of $\{10,20,40,80\}$. The aggregated results averaged over the MuJoCo domains are shown in Fig. 5. Note that we also report the performance of the best baseline. As expected, SUNG with TD3+BC can accommodate preference for both high uncertainty and high value, whereas SUNG with CQL can only accommodate preference for high uncertainty. Furthermore, we remark that SUNG consistently outperforms the best baseline with any choice of $k$, which further verifies the superiority of SUNG.
V-E2 Analysis of Adaptive Exploitation
Moreover, we explore the choice of percent $p$ of identified OOD samples from the mini-batch in adaptive exploitation. We tune it in the range of $\{1,5,10,20,50\}$. The aggregated results averaged over MuJoCo domains are shown in Fig. 6. Predictably, setting $p$ too conservatively or too aggressively may adversely degrade the performance. We find that $p=5$ performs the best when combined with either TD3+BC or CQL, achieving the desired trade-off between performance and stability. Furthermore, we want to highlight that SUNG outperforms the best baseline with most choices of $p$, which further emphasizes the superiority of SUNG.
V-F SUNG with Other Advanced Techniques (RQ4)